9.) What is the molarity of a solution that is 12.5% by mass calcium chloride in water?
Answer #1

Let density of solution be same as density of water which is 1 g/mL

Let volume of solution be 1 L

volume , V = 1 L

= 1*10^3 mL

density, d = 1.0 g/mL

use:

mass = density * volume

= 1 g/mL *1*10^3 mL

= 1*10^3 g

This is mass of solution

mass of CaCl2 = 12.5 % of mass of solution

= 12.5*1000.0/100

= 125.0 g

Molar mass of CaCl2,

MM = 1*MM(Ca) + 2*MM(Cl)

= 1*40.08 + 2*35.45

= 110.98 g/mol

mass(CaCl2)= 125.0 g

use:

number of mol of CaCl2,

n = mass of CaCl2/molar mass of CaCl2

=(1.25*10^2 g)/(1.11*10^2 g/mol)

= 1.126 mol

volume , V = 1 L

use:

Molarity,

M = number of mol / volume in L

= 1.126/1

= 1.126 M

Answer: 1.13 M